# Theoretical analysis of performance limitation of computational refocusing in optical coherence tomography

### v1

@@ 3 Criteria for maximum correctable defocus
Two factors limit the maximum correctable amount of defocus: the lateral image sampling density and the confocality. The former factors both the point-scanning OCT and spatially-coherent FFOCT, whereas the latter factor only affects point-scanning OCT. The maximum correctable defocus amounts defined by these factors are derived in the following sections.
@@ 3.1.1 Nyquist criterion for point-scanning OCT
The PSF of point-scanning OCT [Eq. (8)]is given by the product of the real Gaussian amplitude [Eq. (6)] and the phase-only function $\exp\left\{i{\varphi_{\mathrm{ps}}}({\boldsymbol{r}};z_{d})\right\}$, where ${\varphi_{\mathrm{ps}}}$ is the phase defined in Eq. (7). Therefore, the spatial frequency spectrum of the PSF is given by convolution of the Fourier transforms of the real Gaussian amplitude and the phase-only function. As the defocus increases, the real Gaussian amplitude becomes broader, and thus, its spatial frequency spectrum becomes narrower. On the other hand, as the defocus increases, the phase-only function consists of higher frequency components, especially at the periphery (i.e., at larger $r$, where $r=\left|{\boldsymbol{r}}\right|=\sqrt{x^{2}+y^{2}}$). As a result, the Nyquist criterion is governed by the Nyquist frequency of the phase-only function in this case.
The phase-only function is a quadratic function of ${\boldsymbol{r}}$, and its local frequency increases as $r$ increases. To sample the OCT signal to allow it to be refocused, the lateral sampling density should be high enough when compared with the local frequency. For the phase-only function, the Nyquist condition can be described as follows: “the adjacent sampling points should have a phase difference smaller than or equal to $\pi$,”which can be written as:
@@
where we replaced ${\boldsymbol{r}}$ with $x$ without losing generality. Here, $\Delta{\varphi_{\mathrm{ps}}}$ is the phase difference between adjacent sampling points around $x$ and $\Delta x$ is the lateral sampling distance. $x$ is a generalized lateral position which can be in any lateral direction. We also assume $y$ is the counter part of $x$ and is oriented along the direction orthogonal to $x$. The origin of the $(x,y)$ coordinates is collocated with the center of the PSF. Note here that, to derive the final criteria, $x$ should be along the direction in which the pixel separation reaches a maximum. Specifically, if the lateral sampling is isotropic, then $x$ is neither the fast nor slow scan directions, but instead is along a direction at 45 degree to the slow or fast scan directions.
As the equation shows [Eq.(13)], the absolute phase increments linearly increases by $x$, and thus it reaches a maximum at the periphery of the PSF. Here, we can reasonably define the radius of the PSF as the $1/e$-radius of the amplitude, i. e., $w(z_{d})/\sqrt{2}$, and thus, the maximum absolute phase increment, which is observed at the periphery of the PSF, becomes
@@
where $\mathrm{max}(\quad)_{x}$ represents the maximum over $x$, and this maximum is obtained at $x=w(z_{d})/\sqrt{2}$. Note that the variable $x$ is now considered to be a parameter and the $z_{d}$ is newly treated as a variable rather than a parameter in this equation.
Figure 4: The phase increments per mm at the periphery of the PSF, which corresponds to Eq. (15) and Eq. (20) but without absolute operations. The blue and red curves correspond to the cases of point-scanning OCT and spatially-coherent FFOCT, respectively. The horizontal axis corresponds to the normalized defocus distance $\zeta_{d}$, which takes a value of ±$\pm 1$ when the absolute defocus distance is equal to the Rayleigh distance. The phase increments takes its maximum and minimum values when $\zeta_{d}$ approaches to +$\infty$+ ∞ and $\infty$- ∞, respectively. For the figure plot, we assumed a typical specification of objective lens with $NA=0.13$, $\lambda_{0}=840nm$, $n=1.38$ and $1/e$-radius of the PSF amplitude acting as the defocus radius for point-scanning OCT and spatially-coherent FFOCT.
@@
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(\zeta_{d};x)\right|\right)_{x}=\left|\frac{2\sqrt{2}}{nw_{0}}\frac{1}{1+1/\zeta_{d}^{2}}\Delta x\right|.$ (15)
See the Appendix for the detailed derivation of this equation. To aid intuitive understanding, $\Delta{\varphi_{\mathrm{ps}}}(\zeta_{d};x)$ at the PSF periphery (i.e., $x=w(z_{d})/\sqrt{2}$) is plotted as a function of $\zeta_{d}$ in Fig. 4.
@@
$\mathrm{max}\left(\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(\zeta_{d};x)\right|\right)_{x}\right)_{\zeta_{d}}=\lim_{\zeta_{d}\rightarrow\pm\infty}\left|\frac{2\sqrt{2}}{nw_{0}}\frac{1}{1+1/\zeta_{d}^{2}}\Delta x\right|=\frac{2\sqrt{2}}{nw_{0}}\Delta x.$ (16)
Because we took the maximum of a maximum, the right hand side of this equation represents the maximum absolute phase that can occur when we sample the OCT signal with a sampling distance of $\Delta x$.
@@
$\Delta x\leq\frac{\pi nw_{0}}{2\sqrt{2}}.$ (17)
This criterion can be interpreted as follows. Specifically, as long as the longest adjacent-pixel separation is smaller than $\pi n/2$ times the $1/e$-radius of the diffraction-limit PSF amplitude ($w_{0}/\sqrt{2}$), the defocus is correctable via holographic refocusing, regardless of the defocus distance. In other words, as long as the the longest adjacent-pixel separation is smaller than $\pi n/4$ times the in-tissue lateral resolution, i.e., $1/e^{2}$-diameter of the diffraction-limited PSF intensity ($\sqrt{2}w_{0}$), the defocus remains correctable.
When we consider isotropic lateral sampling specifically, this condition can be restated as follows:
$\Delta x^{\prime}\leq\frac{\pi nw_{0}}{4},$ (18)
where $\Delta x^{\prime}$ is the adjacent-pixel separation along either the fast or slow scanning direction. In other words, as long as the lateral pixel distance along the scan directions is smaller than $\pi n/4\sqrt{2}$ times the in-tissue lateral resolution defined by the $1/e^{2}$-width of the PSF intensity ($\sqrt{2}w_{0}$), the defocus is correctable. This condition is also equivalent to $\pi/4\sqrt{2}$ times the in-air lateral resolution defined by $1/e^{2}$-width of the PSF intensity ($\sqrt{2}nw_{0}$).
@@ 3.1.2 Nyquist criterion for spatially-coherent FFOCT
$\left|\Delta{\varphi_{\mathrm{sc}}}(x;z_{d})\right|=\left|\frac{\partial}{\partial x}\left(\frac{nk_{0}x^{2}}{2R(z_{d})}\right)\Delta x\right|=\left|\frac{nk_{0}x}{R(z_{d})}\Delta x\right|.$ (19)
and the maximum absolute phase increment at the normalized defocus depth of $\zeta_{d}$ becomes
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{sc}}}(\zeta_{d};x)\right|\right)_{x}=\left|\frac{2}{nw_{0}}\frac{1}{1+1/\zeta_{d}^{2}}\Delta x\right|.$ (20)
[…]
$\mathrm{max}\left(\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{sc}}}(\zeta_{d};x)\right|\right)_{x}\right)_{\zeta_{d}}=\lim_{\zeta_{d}\rightarrow\pm\infty}\left|\frac{2}{nw_{0}}\frac{1}{1+1/\zeta_{d}^{2}}\Delta x\right|=\frac{2}{nw_{0}}\Delta x.$ (21)
To fulfill the Nyquist condition, the value must be smaller than or equal to $\pi$, and this gives us the Nyquist criterion for holographic refocusing for spatially-coherent FFOCT, as follows:
$\Delta x\leq\frac{\pi nw_{0}}{2}.$ (22)
Specifically, for isotropic lateral sampling, the criterion becomes:
$\Delta x^{\prime}\leq\frac{\pi nw_{0}}{2\sqrt{2}},$ (23)
where $x^{\prime}$ is the pixel separation along either the fast or slow scanning direction.
These Nyquist criteria suggest that, regardless of the defocus distance, the defocus is correctable as long as the horizontal or vertical adjacent pixel separation remains smaller than or equal to $\pi n/4\sqrt{2}$ times the $1/e^{2}$-width of the in-tissue diffraction-limit PSF intensity ($2w_{0}$). Additionally, the criterion is equivalent to $\pi/4\sqrt{2}$ times the $1/e^{2}$-width of the in-air diffraction-limit PSF intensity ($2nw_{0}$).
It should be noted here that spatially-coherent FFOCT does not have a confocal pinhole and is thus free from the confocality limit that will be discussed in the next section. Therefore, this criterion is only the requirement to ensure that the defocus is correctable for spatially-coherent FFOCT.
@@ 3.2 Confocality-limit criterion for point-scanning OCT
We assume that the total signal intensity (${\mathrm{I}_{\mathrm{ps}}}$) captured at a specific depth in the point-scanning OCT is proportional to a confocal function $h(z_{d})$. The confocal function is defined as an intensity integral of PSF for point-scanning OCT over the lateral integration direction, as follows:
[…]
where $\mathrm{PSF}_{\mathrm{ps}}(r)$ is the PSF defined by using Eq. (8) along with the substitution of ${\boldsymbol{r}}\cdot{\boldsymbol{r}}=r^{2}$. The integration is conducted over the radial direction and $r$ after the squared PSF represents the Jacobian. Here, we did not take the light attenuation by scattering from the sample into account for simplicity. The details of this issue will be discussed in Section 5.3.
The dB-scaled intensity profile of Eq. (24) is shown as a function of the normalized defocus distance $\zeta_{d}=z_{d}/z_{R}$ in Fig. 5, where the peak at the in-focus depth is set as 0 dB because
$10\log_{10}{\mathrm{I}_{\mathrm{ps}}}^{\prime}(\zeta_{d})=10\log_{10}\frac{1}{1+\zeta_{d}^{2}},$ (25)
where ${\mathrm{I}_{\mathrm{ps}}}^{\prime}(\zeta_{d})$ is the total signal intensity when normalized with respect to its maximum $w_{0}^{2}/8$.
Figure 5: Intensity profile of the peak intensity of the refocused signal when computed using Eq. (25). The yellow dashed line and the red dashed lines indicate the noise level and cut-off defocus distances, respectively. The signal becomes observable after refocusing if the defocus distances are less than the cut-off defocus distances (as indicated by the double-sided red arrow). Here is an example for SNR of 20 dB, the critical defocus distance becomes $10z_{R}$. For the figure plot, we assume the same specification as Fig. 4.
This intensity profile can be regarded as the peak intensity profile of refocused signal. By assuming a specific SNR, we can then find the critical defocus distance $z_{c}$ at which the SNR becomes 0 dB and the signal disappears using
@@
where $\zeta_{c}=z_{c}/z_{R}$ is the normalized critical defocus and $\mathrm{SNR_{dB}}$ is the SNR in dB scale. This definition of the critical defocus, i.e., the confocality limit criterion, can be rewritten as:
@@
This definition of confocality limit criterion is also illustrated schematically in Fig. 5. Here, the dashed yellow line represents the noise level and the red dashed lines indicate the critical defocus distance.
@@ 3.3 Summary for maximum sampling distance
For point-scanning OCT, the maximum correctable defocus is dominated by the criterion that is more sever between the Nyquist criterion and the confocality-limit criterion. According the Nyquist criterion, the defocus remains correctable as long as the pixel separation along the fast and slow scan direction is smaller than $\pi n/4\sqrt{2}$ times the diffraction limit of the in-tissue lateral resolution defined as $1/e^{2}$-width of the PSF intensity, or smaller than $\pi/4\sqrt{2}$ times the corresponding in-air lateral resolution, regardless of the defocus distance. According to the confocality limit criterion, the defocus distance should be smaller than the critical defocus distance $z_{c}$ defined in Eq. (27); otherwise the signal cannot be observed, even after holographic refocusing.
For spatially-coherent FFOCT, the Nyquist criterion is only one criterion. According to this criterion, the defocus is correctable as long as the horizontal and vertical pixel separation is smaller than $\pi n/4\sqrt{2}$ times the in-tissue lateral diffraction-limit resolution, or smaller than $\pi/4\sqrt{2}$ times the corresponding in-air lateral resolution. Notethat the Nyquist criteria for both types of OCT does not depend on the defocus distance. In addition, if we assume the same optical systems parameters are used for both point-scanning OCT and spatially-coherent FFOCT, the sampling density requirement of spatially-coherent FFOCT is $\sqrt{2}$ times looser than that for point-scanning OCT. Please note that, in this case, the lateral resolution for spatially-coherent FFOCT is $\sqrt{2}$ times lower than that for point-scanning OCT. In other words, if design the same resolution for point-scanning and spatially-coherent FFOCT, the Nyquist criterion is the same for two systems, but the incident beam width of spatially-coherent FFOCT is smaller than that of point-scanning OCT. This might be a system design advantages for spatially-coherent FFOCT.
Furthermore, unlike point-scanning OCT, FFOCT does not employ confocal gating, meaning that out-of-focus light is still captured and can be computationally refocused. And thus, the absence of the confocality make spatially-coherent FFOCT advantageous for computational refocusing.
@@ 4 Examples cases
The maximum correctable defocus and the related system specifications have been analyzed for several OCT systems, with results as summarized in Table Table1. A Jones-matrix swept-source OCT (JM-SSOCT) system constructed by the authors [21] was included as a representative of scanning swept-source OCT with relatively low lateral resolution (approximately 18 $\mu$m ). Please note that the JM-SSOCT defines the in-air lateral resolution as $1/e^{2}$ of the beam spot diameter. The beam spot diameter corrsponds to the $1/e^{2}$-width of the PSF amplitude. In this study, the in-air lateral resolution is defined as the $1/e^{2}$-width of the PSF intensity ($\sqrt{2}nw_{0}$). This definition leads to the in-air lateral resolution of $18/\sqrt{2}$ $\mu$m (12.73 $\mu$m ) in this manuscript. Although this system is polarization sensitive, it does not affect our analyses. A standard spectral-domain OCT (SD-OCT) system operating in the 840 nm band that was also built by the authors [22, 23] was included as a representative example of a relatively high-resolution scanning OCT system(around 4.9 $\mu$m ). This corresponds to a resolution of $4.9/\sqrt{2}$ $\mu$m (3.50 $\mu$m ) in this manuscript. This system has been used widely as a basis for holographic signal processing studies, including computational refocusing applications[24, 25].
For spatially-coherent FFOCT, we included our own spatially-coherent FFOCT system[26] because this system follows a standard spatially-coherent FFOCT and all information about the system’s design and specifications is available. One important variation of spatially-coherent FFOCT is off-axis spatially-coherent FFOCT. Time-domain [27] and swept-source [28] off-axis spatially-coherent FFOCT are included in our analysis. The literature entry for the latter system also provided the parameters for the on-axis spatially-coherent FFOCT system, which may have been used to perform frequency phase-mask-based aberration correction [29]. We also included spatio-temporal-optical-coherence tomography (STOC-T) system demonstrated by Wojtkowski et al.[30]. STOC-T can be regarded as a variation of spatially-coherent FFOCT that includes a wavefront modulation mechanism. The specifications of the systems are partially available in the literature[31], and the STOC-T system has also been used for holographic refocusing [32, 33]. The in-air lateral resolution was unified by $1/e^{2}$-width of the PSF intensity ($2nw_{0}$).
@@
Note that the Nyquist criterion for both point-scanning OCT and spatially-coherent FFOCT can be summarized as, “if the pixel separation/lateral resolution (i.e., fractional pixel separation) is smaller than or equal to 55.5%, the defocus can be corrected regardless of the original defocus amount.” In the Jones-matrix SS-OCT with the a configuration of 512 $\times$ 512 lateral pixels with a 3 mm $\times$ 3 mm or smaller field of view, this criterion is fulfilled. The scanning SD-OCT has higher lateral resolution than the Jones-matrix SS-OCT, and thus it only a small field of view, e. g., 1 mm $\times$ 1 mm, which is almost compatible with this criterion as long as the lateral pixel number is 512 $\times$ 512 pixels. The on-axis SS-FFOCT of the University of Tsukuba was designed to be compatible with this criterion, with a 33.3% fraction pixel separation. However, for the STOC-T, further computational refocusing improvements can be expected by increasing the pixel density. Although the off-axis SS-FFOCT of the University of Lübeck fulfills the criterion, we also need to consider the off-axis nature of this system, as we will discuss in Section 5.1.
@@
Table 1: Specifications and critical defocus distances for example OCT systems. Pix., res., sep., and FOV are abbreviations for pixel, resolution, separation, and field of view, respectively. The critical defocus distance is the confocality limit-based critical defocus distance that was defined in Section 3.2.

### v3

@@ 3 Criteria for maximum correctable defocus
Two factors limit the maximum correctable amount of defocus: the lateral image sampling density and the confocality. The former factor both the point-scanning OCT and spatially-coherent FFOCT, whereas the latter factor only affects point-scanning OCT. The maximum correctable defocus amounts defined by these factors are derived in the following sections.
@@ 3.1.1 Nyquist criterion for point-scanning OCT
The PSF of point-scanning OCT [Eq. (8)] is given by the product of the real Gaussian amplitude [Eq. (6)] and the phase-only function $\exp\left[i{\varphi_{\mathrm{ps}}}({\boldsymbol{r}};z_{d})\right]$, where $\varphi_{\mathrm{ps}}$} is the phase defined in Eq. (7). Therefore, the spatial frequency spectrum of the PSF is given by convolution of the Fourier transform of the real Gaussian amplitude and the Fourier transform of the phase-only function. As the defocus increases, the real Gaussian amplitude becomes broader, and thus, its spatial frequency spectrum becomes narrower. On the other hand, as the defocus increases, the phase-only function consists of higher frequency components, especially at the periphery (i.e., at larger $r$, where $r=\left|{\boldsymbol{r}}\right|=\sqrt{x^{2}+y^{2}}$). As a result, the Nyquist criterion is governed by the Nyquist frequency of the phase-only function in this case.
The phase-only function is a quadratic function of $\boldsymbol{r}$}, and its local frequency increases as $r$ increases. To sample the OCT signal to allow it to be refocused, the lateral sampling density should be high enough when compared with the local frequency. For the phase-only function, the Nyquist condition can be described as follows: “the adjacent sampling points should have a phase difference smaller than or equal to $\pi$,” which can be written as
@@
where we replaced $\boldsymbol{r}$} with $x$ without losing generality. Here, $\Delta{\varphi_{\mathrm{ps}}}$ is the phase difference between adjacent sampling points around $x$ and $\Delta x$ is the lateral sampling distance. $x$ is a generalized lateral position which can be in any lateral direction. We also assume $y$ is the counter part of $x$ and is oriented along the direction orthogonal to $x$. The origin of the $(x,y)$ coordinates is collocated with the center of the PSF.
As Eq. (13) shows, the absolute phase increments linearly increases by $x$, and thus it reaches a maximum at the periphery of the PSF. Here, we can reasonably define the radius of the PSF as the $1/e$-radius of the amplitude, i. e., $w(z_{d})/\sqrt{2}$, and thus, the maximum absolute phase increment, which is observed at the periphery of the PSF, becomes
@@
where $\mathrm{max}(\quad)_{x}$ represents the maximum over $x$, and this maximum is obtained at $x=w(z_{d})/\sqrt{2}$. Now we consider the $x$ as a parameter with a certain value, while $z_{d}$ will be treated as a variable rather than a parameter.
[…]
@@
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(\zeta_{d};x)\right|\right)_{x}=\left|\frac{2\sqrt{2}}{w_{0}}\frac{\operatorname{sgn}(\zeta_{d})}{\sqrt{1+1/\zeta_{d}^{2}}}\Delta x\right|.$ (15)
See the Appendix A for the detailed derivation of this equation. To aid intuitive understanding, $\Delta{\varphi_{\mathrm{ps}}}(\zeta_{d};x)$ at the PSF periphery (i.e., $x=w(z_{d})/\sqrt{2}$) is plotted as a function of $\zeta_{d}$ in Fig. 4.
@@
$\mathrm{max}\left(\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{ps}}}(\zeta_{d};x)\right|\right)_{x}\right)_{\zeta_{d}}=\lim_{\zeta_{d}\rightarrow\pm\infty}\left|\frac{2\sqrt{2}}{w_{0}}\frac{\operatorname{sgn}(\zeta_{d})}{\sqrt{1+1/\zeta_{d}^{2}}}\Delta x\right|=\frac{2\sqrt{2}}{w_{0}}\Delta x.$ (16)
Because we took the maximum of the maximum, the right hand side of this equation represents the maximum absolute phase increment that can occur when we sample the OCT signal with a sampling distance of $\Delta x$.
@@
$\Delta x\leq\frac{\pi w_{0}}{2\sqrt{2}}.$ (17)
This criterion can be interpreted as follows. Specifically, as long as the adjacent-pixel separation is smaller than $\pi/2$ times the $1/e$-radius of the diffraction-limit PSF amplitude ($w_{0}/\sqrt{2}$), the defocus is correctable via holographic refocusing, regardless of the defocus distance. In other words, as long as the adjacent-pixel separation is smaller than $\pi/4$ times the lateral resolution, i.e., $1/e^{2}$-diameter of the diffraction-limited PSF intensity ($\sqrt{2}w_{0}$), the defocus remains correctable.
@@ 3.1.2 Nyquist criterion for spatially-coherent FFOCT
$\left|\Delta{\varphi_{\mathrm{sc}}}(x;z_{d})\right|=\left|\frac{\partial}{\partial x}\left(\frac{nk_{0}x^{2}}{2R(z_{d})}\right)\Delta x\right|=\left|\frac{nk_{0}x}{R(z_{d})}\Delta x\right|,$ (18)
and the maximum absolute phase increment at the normalized defocus depth of $\zeta_{d}$ is
$\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{sc}}}(\zeta_{d};x)\right|\right)_{x}=\left|\frac{2}{w_{0}}\frac{\operatorname{sgn}(\zeta_{d})}{\sqrt{1+1/\zeta_{d}^{2}}}\Delta x\right|.$ (19)
The non-absolute version of this equation is plotted in Fig. 4 (orange curve). Is it can be seen in the plot, the maximum absolute phase increment approaches its maxima asymptotically as $\zeta_{d}$ approaches $\pm\infty$ as
$\mathrm{max}\left(\mathrm{max}\left(\left|\Delta{\varphi_{\mathrm{sc}}}(\zeta_{d};x)\right|\right)_{x}\right)_{\zeta_{d}}=\lim_{\zeta_{d}\rightarrow\pm\infty}\left|\frac{2}{w_{0}}\frac{\operatorname{sgn}(\zeta_{d})}{\sqrt{1+1/\zeta_{d}^{2}}}\Delta x\right|=\frac{2}{w_{0}}\Delta x.$ (20)
To fulfill the Nyquist condition, the value must be smaller than or equal to $\pi$, and this gives us the Nyquist criterion for holographic refocusing for spatially-coherent FFOCT, as follows
$\Delta x\leq\frac{\pi w_{0}}{2}.$ (21)
These Nyquist criteria suggest that, regardless of the defocus distance, the defocus is correctable as long as the adjacent pixel separation remains smaller than or equal to $\pi/4$ times the $1/e^{2}$-width of the diffraction-limit PSF intensity ($2w_{0}$).
It should be noted here that spatially-coherent FFOCT does not have a confocal pinhole and is thus free from the confocality limit that will be discussed in the next section. Therefore, this Nyquist criterion is only the requirement to ensure that the defocus is correctable for spatially-coherent FFOCT.
@@ 3.2 Confocality-limit criterion for point-scanning OCT
We assume that the total signal intensity ($\mathrm{I}_{\mathrm{ps}}$}) captured at a specific depth in the point-scanning OCT is proportional to a confocal function $h(z_{d})$. The confocal function is defined as an intensity integral of PSF for point-scanning OCT over the lateral integration direction, as follows
${\mathrm{I}_{\mathrm{ps}}}(z_{d})\propto h(z_{d})\propto\int_{0}^{2\pi}\int_{0}^{+\infty}\left|\mathrm{PSF}_{\mathrm{ps}}(r)\right|^{2}r\,\mathrm{d}r\,\mathrm{d}\theta=\frac{{\pi w_{0}}^{2}}{4[1+(z_{d}/z_{R})^{2}]},$ (22)
where $\mathrm{PSF}_{\mathrm{ps}}(r)$ is the PSF defined by using Eq. (8) along with the substitution of ${\boldsymbol{r}}\cdot{\boldsymbol{r}}=r^{2}$. Here, we did not take the light attenuation by scattering from the sample into account for simplicity. The details of this issue will be discussed in Section 5.2.
Figure 5: Intensity profile of the peak intensity of the refocused signal. The orange line and the red dashed line indicate the noise level and the critical defocus distance, respectively. The signal becomes observable after refocusing if the defocus distance is less than the critical defocus distance.
The dB-scaled profile of Eq. (22) is shown as a function of the normalized defocus distance $\zeta_{d}=z_{d}/z_{R}$ in Fig. 5, where the peak at the in-focus depth is set as 0 dB. This profile can be regarded as the peak intensity profile of refocused signal. By assuming a specific SNR, we can then find the critical defocus distance $z_{c}$ at which the SNR becomes 0 dB and the signal disappears using
@@
where $\zeta_{c}=z_{c}/z_{R}$ is the normalized critical defocus and $\mathrm{SNR_{dB}}$ is the SNR in dB scale. This definition of the critical defocus, i.e., the confocality-limit criterion, can be rewritten as
@@
This definition of confocality-limit criterion is also illustrated schematically in Fig. 5. Here, the dashed yellow line represents the noise level and the red dashed lines indicate the critical defocus distance.
@@ 3.3 Summary for maximum sampling distance
For point-scanning OCT, the maximum correctable defocus is limited by the more stringent of two criteria; the Nyquist criterion and the confocality-limit criterion. According the Nyquist criterion, the defocus remains correctable as long as the pixel separation is smaller than $\pi/4$ times the diffraction limit of the lateral resolution defined as $1/e^{2}$-width of the PSF intensity regardless of the defocus distance. According to the confocality-limit criterion, the defocus distance should be smaller than the critical defocus distance $z_{c}$ defined in Eq. (24), otherwise the signal cannot be observed, even after holographic refocusing.
For spatially-coherent FFOCT, the Nyquist criterion is only one criterion. The final description of this criterion becomes the same with that of the point-scanning OCT. Namely, the defocus is correctable as long as the horizontal and vertical pixel separation is smaller than $\pi/4$ times the lateral diffraction-limit resolution . Note that the Nyquist criteria for both types of OCT does not depend on the defocus distance. Unlike point-scanning OCT, spatially-coherent FFOCT does not employ confocal gating, meaning that out-of-focus light is still captured and can be computationally refocused. The absence of the confocality make spatially-coherent FFOCT advantageous for computational refocusing.
@@ 4 Examples cases
[…]
These two systems have been used for the studies of holographic signal processing and computational refocusing [37, 39, 40].
As an example of spatially-coherent FFOCT system, we included our own spatially-coherent FFOCT system[41] because this system follows a standard spatially-coherent FFOCT configuration and all system design parameters and specifications are available. One important variation of spatially-coherent FFOCT is off-axis spatially-coherent FFOCT. And hence, time-domain [42] and swept-source [43] off-axis spatially-coherent FFOCT are included in our analysis.
We also include spatio-temporal-optical-coherence tomography (STOC-T) demonstrated by Auksorius et al.[44, 45]. The STOC-T system has also been used for holographic aberration correction [46]. Since STOC-T can be regarded as a spatially-coherent FFOCT when the wavefront modulation is turned off, we included their optical parameters in the table for assessing the optical configuration.
In Table 1, the lateral resolutions of the point-scanning OCT systems were shown as the $1/e^{2}$-width of the PSF intensity ($2w_{0}$) . In the case of spatially-coherent FFOCT, the lateral resolution was calculated with the formula $4\times 0.22\frac{\lambda}{\mathrm{NA}_{\mathrm{col}}}$ according to the Gaussian profile fitting to the Airy disk pattern (Eq. (21) in Ref. [35]), where NAcol is the collection NA.
@@
Note that the Nyquist criterion for both point-scanning OCT and spatially-coherent FFOCT can be summarized as, “if the ratio of pixel separation to lateral resolution (i.e., fractional pixel separation) is smaller than or equal to 78.5% (i.e., $\pi/4$), the defocus can be corrected regardless of the original defocus amount.” In the Jones-matrix SS-OCT with the a configuration of 512 $\times$ 512 lateral pixels with a 3 mm $\times$ 3 mm or smaller field of view, this criterion is fulfilled. The scanning SD-OCT has higher lateral resolution than the Jones-matrix SS-OCT, and thus it requires smaller field of view, such as 1 mm $\times$ 1 mm, as long as far as we kept the same lateral pixel number; 512 $\times$ 512 pixels. The on-axis SS-FFOCT of the University of Tsukuba was designed to sufficiently fulfill this criterion. The STOC-T systems have a fractional pixel separation fulfilling the criteria. The off-axis SS-FFOCT of the University of Lübeck also fulfills the criterion, but the criterion becomes tighter in practice for the off-axis configuration as we will discuss in Section 5.1.
@@
Table 1: Specifications and critical defocus distances for example OCT systems. Pix., res., sep., and FOV are abbreviations for pixel, resolution, separation, and field of view, respectively. The critical defocus distance is the confocality-limit-based critical defocus distance that was defined in Section 3.2.